# On A Scale From 1 to 5: Quantifying Hallucination in Faithfulness Evaluation

### v2

@@ A.3 Additional Results
We also experimented on Qwen-7B & 72B Yang et al. (2024), but did not include the results in the main paper because Qwen-7B occasionally (about 10% of the times) responded with Chinese reasoning, which caused difficulties in the comparison and analysis. Furthermore, we quickly explored fine-tuning HHEM with intrinsic and extrinsic combined synthetic data, however, the initial results indicate that the HHEM-combined was a little too strict in scoring gold data. This was likely due to the hallucination samples being double in size than the gold samples after combining the datasets. Due to resource limitation, we did not further adjust the training data to achieve desired results. In Table 9, we show the residual values by task for Qwen, HHEM-combined, and all models presented in the main paper as reference.
@@ A.4 Additional Discussion on Evaluation Subjectivity
One may argue that a rubric-based approach is subject to to the particular use case which can be challenging across different contexts. This concern stands for not only the method proposed in this paper, but also other polarity-based faithfulness classification models. Similar to any traditional machine learning methods, the definition or gold data is defined by human annotators. A model can also be over-fitting to a particular use case if the gold labels are very domain specific. For example, in opinion mining, "A is very fast" can indicate positive opinion if "A" is a calculator, but it can also indicate negative opinion if "A" is a person driving in the resident area full of kids. Now the example might seem a bit extreme, the point we’re trying to make is that while consistency is very important, subjectivity is not 100% bad and it can sometimes help domain-specific adaptions.
In our case, the subjectivity only resides in the quantifiable metrics, but not in the domain itself (e.g. travel). In Table 5, we attached the prompt used to score summaries during the experiments. The prompt is domain independent which can be applied to any summarization tasks. Because some quantifiable metrics such as adjective usage and tone of voice can be subjective depending on the use case, having the ability for the user to tailor the rubrics will best help the LLM judge fit to that particular use case. For example, it is possible in stylized text generation that having LLMs add descriptive adjectives and verbs is desired in the generation, in which case, the user can adjust the “adjective consistency” requirements in the rubrics to fit that use case. We intended to introduce an adaptable solution, and similar to any other ML models, small adjustment might still be required when adapting to granular details.
@@
Due to the lack of organic hallucinations in applications, we focused on synthetic hallucinations generated based on real-world observations. One may raise the concern that it is risky to draw conclusions on synthetic hallucinations over organic ones, as the distribution may not be fully modeled. We agree that organic hallucinations are more valuable and that in ideal scenarios, we should develop such faithfulness models using orgnic data as much as we can. However, we believe in the real-world case by case scenario, it is rather difficult to obtain a large portion of organic hallucinations for development. We believe for a well-defined system with a trained developer, even the initial development stage would suffer from insufficient negative examples. However, this does not mean the evaluation should be omitted. It is rather common for developer to generate synthetic test cases (or edge cases) for testing. Thus, we tailored the synthetic data generation for each use case to tackle tricky and subtle inconsistencies.
We believe the conclusion on the experiments can still be generalized because 1) Real-world dataset often contains more noise and are more complex in nature than academic gold dataset. For example, ConvoTS datasets contains many turns of redundant user interactions due to the limitation of the virtual assistant. 2) A portion of the synthetic changes are tricky enough to mimic organic hallucination (see examples in Table 7). Furthermore, even the best LLM judge (GPT-4) failed on some of those tricky cases.
@@
Table 9: Overall residual (absolute error) by task on all models tested. The numeric values match with the visualizations presented in Figures 2. Additional models, Qwen 2.5 7B & 72B and HHEM-combined are included for reference.

### v3

@@ A.3 Additional Results
We also experimented on Qwen-7B & 72B Yang et al. (2024), but did not include the results in the main paper because Qwen-7B occasionally (about 10% of the times) responded with Chinese reasoning, which caused difficulties in comparison and analysis. Furthermore, we quickly explored fine-tuning HHEM with intrinsic and extrinsic combined synthetic data; however, the initial results indicate that the HHEM-combined was a little too strict in scoring gold data. This was likely due to the hallucination samples being double in size than the gold samples after combining the datasets. Due to resource limitation, we did not further adjust the training data to achieve the desired results. In Table 10, we show the residual values by task for Qwen, HHEM-combined, and all models presented in the main paper as a reference.
@@ A.4 Additional Discussion on Evaluation Subjectivity
One may argue that a rubric-based approach is subject to the particular use case, which can be challenging across different contexts. This concern stands for not only the method proposed in this paper, but also other polarity-based faithfulness classification models. Similarly to any traditional machine learning methods, the definition or gold data is defined by human annotators. A model can also be overfitting to a particular use case if the gold labels are very domain-specific. For example, in opinion mining, "Object A is very fast" can indicate a positive opinion if "Object A" is a calculator, but it can also indicate a negative opinion if "Object A" is a car running at full speed in a residential area with young children. Now this example might seem a bit extreme; the core is that while consistency is indeed important, subjectivity is not 100% bad and it can sometimes help domain-specific adaptions.
On the other hand, in our case, the subjectivity only resides in the quantifiable metrics, but not in the domain itself (e.g. travel). We argue that quantifiable metrics should be viewed as a separate component as the domain. We view quantifiable metrics as downstream tasks which can be combined with flexibilities. In Table 6, we attached the prompt used to score summaries during the experiments. The prompt is domain independent, which can be applied to any summarization tasks. Because some quantifiable metrics such as adjective usage and tone of voice can be subjective depending on the use case, having the ability for the user to tailor the rubrics will best help the LLM judge fit to that particular use case. For example, it is possible in stylized text generation that having LLMs add descriptive adjectives and verbs is desired in the generation, in which case the user can adjust the “adjective consistency” requirements in the rubrics to fit that use case. We intended to introduce an adaptable solution, and similar to any other ML models, small adjustment might still be required when adapting to granular details.
Finally, on concerns over generalization. We believe that in the current era of LLMs, generalization of GenAI downstream tasks should not be strictly defined as a "universal model that powers all use cases". Rather, generalization should aim for a "framework" or "template" that can be adopted easily by other domains. With a newer and better model evolving at a monthly basis, we believe the prompt templates and experiment designs we proposed in this paper can be generalized to changes.
@@
Due to the lack of organic hallucinations in applications, we focused on synthetic hallucinations generated based on real-world observations. One may raise the concern that it is risky to draw conclusions on synthetic hallucinations over organic ones, as the distribution may not be fully modeled. We agree that organic hallucinations are more valuable and that in ideal scenarios, we should develop such faithfulness models using organic data as much as we can. However, we believe that in the real-world case-by-case scenario, it is rather difficult to obtain a large portion of organic hallucinations for development. We believe that for a well-defined system with a trained developer, even the initial development stage would suffer from insufficient negative examples. However, this does not mean that the evaluation should be omitted. It is rather common for developers to generate synthetic test cases (or edge cases) for testing. Thus, we tailored the synthetic data generation for each use case to tackle tricky and subtle inconsistencies.
We believe that the conclusion on the experiments can still be generalized because 1) the real-world dataset often contains more noise and is more complex in nature than the academic gold dataset. For example, ConvoTS datasets contain many turns of redundant user interactions due to the limitation of the virtual assistant. 2) A portion of the synthetic changes are tricky enough to mimic organic hallucinations (see examples in Table 8). Furthermore, even the best LLM judge (GPT-4) failed in some of those tricky cases.
@@
Table 10: Overall residual (absolute error) by task on all models tested. The numeric values match with the visualizations presented in Figures 2. For models reported in the main paper, GPT-4 and HHEM-extrinsic performed the best overall. Although Llama3.1-8B and Claude3-Haiku showed very low error rate on half of the tasks, the models are either too strict or too lenient to cover all cases. Additional models, Qwen 2.5 7B & 72B and HHEM-combined are included for reference.